OpenStudy (anonymous):

i will meadel and fan if.... you show me how you get it and explain there is a monthly fee of 35$ at your fitness center. you have spent 700 dollers in fees at the fitness center since you have jioned. write and solve an equation to determine how many months you were a member?

OpenStudy (igreen):

Monthly fee of $35 Spent a total of 700 So what do you think we can use as an equation?

OpenStudy (anonymous):

35x=700

OpenStudy (igreen):

Since you want to know how many months, we can use 'x' to represent months. '35' represents the fee you have to pay per month, so we have: 35x

OpenStudy (igreen):

Yes, that's correct!

OpenStudy (unanimoose):

Wow, that was fast.

OpenStudy (igreen):

Now solve it..divide 35 to both sides.

OpenStudy (anonymous):

can u show me how to do that

OpenStudy (unanimoose):

\[\frac{35x}{35} = \frac{700}{35}\]

OpenStudy (igreen):

\(35x = 700\) \(\dfrac{35x}{35} = \dfrac{700}{35}\) The 35's on the left cancel out: \(\dfrac{\cancel{35}x}{\cancel{35}} = \dfrac{700}{35}\) Leaving us with: \(x = \dfrac{700}{35}\)

OpenStudy (anonymous):

can u guys help me with my question?

OpenStudy (anonymous):

it will equal 20 right?

OpenStudy (igreen):

Yep, you got it, so 20 months